# Fast Algorithms for Fourier extension based on boundary interval data thanks: The research is partially supported by National Natural Science Foundation of China (No. 12171455, RSF-NSFC 23-41-00002)

### v4

@@ 5.2 Testing the parameter $m_{\Delta}$
$\gamma_{\Delta}$ $8$ $4$ $2$ $1$ $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$T_{\Delta}$ $1.1$ $1.2$ $2.3$ $6$ $12$ $24$ $50$
@@ 6.1 Testing and comparison of algorithm performance
Based on the previous tests, we have provided specific values for the parameters involved in the algorithm. In this subsection, the performance of the proposed algorithm is tested. All tests were computed on a Windows 10 system with 16 GB of memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b. We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in [5], the Lyons algorithm (M3) in [12] and the implicit projection algorithm (M4) in [14]. For convenience of numerical implementation, for M3, we set parameters $T=2,\gamma=2$. For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$, $T=2,\gamma=2$, and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different. We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve machine precision).
@@
Fig. 12 shows the numerical results for the first case. As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but M2 cannot achieve superalgebraic convergence using polynomial approximation, so the value of $M$ required is significantly larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1. At this time, the number of nodes $M$ required by the four algorithms to achieve machine accuracy was relatively small, therefore, there was no obvious difference in performance.
@@
In the third case, because the function oscillates near the boundary, the algorithm in this paper cannot avoid this singularity. For functions $f_{7}$ and $f_{8}$, the maximum frequency of the function appears in the boundary interval. According to (27), the number of nodes required by the proposed algorithm is $T_{\Delta}\gamma_{\Delta}=6$ times the maximum frequency, whereas the number of nodes required by M3 and M4 is $T\gamma=4$ times the maximum frequency. This is consistent with the results presented in Figures 2 and 3. For $f_{9}$, because the internal frequency is higher, the number of nodes required by the proposed algorithm is less than that of M3 and M4. As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve similar accuracy.
@@ 6.2 Algorithm improvement: boundary grid refinement
$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$ (38)
@@
$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~{}J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},~{}J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$ (39)
@@
The extension function $g^{R}_{c}(x^{R}_{j}),j=1,2,\ldots,L^{R}_{\Delta}$ was also obtained according to the method described in section 3. And then we obtain a function $f^{R}_{c}$ with a period of $2+\lambda$, where $\lambda$ is same as in (9). The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $[-1,1+\lambda)$ are
@@
Figure 15: The approximation error against $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
@@
Figure 16: Comparisons of algorithms M1 and M4.
Now we take $\gamma=1,T_{\Delta}=6,m_{\Delta}=25$ to test the influence of parameter $R$ on the algorithm. As can be seen from Fig. 15, when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part. When $R\leq 4$, it can be observed that the refinement of grid near the boundary improves the approximation error. When $\omega=50$, the improvement is not obvious, because the number of nodes $M=500>T_{\Delta}\gamma_{\Delta}\omega$ has exceeded the threshold for resolving this frequency function. Therefore, when facing high-frequency functions, we recommend using parameters
@@
for calculation.
@@ Acknowledgments
The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes in [5] and [12] to facilitate comparative analysis in this work.

### v5

@@ 5.2 Testing the parameter $m_{\Delta}$
$\gamma_{\Delta}$ 88 44 22 11 $\frac{1}{2}$ $\frac{1}{4}$ $\frac{1}{8}$
$T_{\Delta}$ $1.1$ $1.2$ $2.3$ 66 $12$ $24$ $50$
@@ 6.1 Testing and comparison of algorithm performance
Based on the previous tests, we have provided specific values for the parameters involved in the algorithm. In this subsection, the performance of the proposed algorithm is tested. All tests were computed on a Windows 10 system with 16 GB of memory, Intel(R) Core(TM)i7-8500U CPU1.80GHz using MATLAB 2016b. We compared the boundary interval algorithm (M1) proposed in this paper with the FC-Gram algorithm (M2) in bruno2010high , the Lyons algorithm (M3) in Lyon2011 and the implicit projection algorithm (M4) in Matthysen2016FAST . For M2, we use the codes at https://github.com/oscarbruno/FC and take the parameters $d=13$, $C=33$. For the specific meaning of the parameters, please refer to 2016An . We also tested other parameter values. As $d$ increases, the algorithm becomes unstable. A large $d$ value (such as $d=30$) will prevent the algorithm from reaching near machine precision. According to our test results, if our approximation goal is to achieve near machine accuracy when $M$ is relatively small, then the optimal value of $d$ is approximately $12\sim 15$. For M3, we set parameters $T=2,\gamma=2$. For M4, we selected the best result from the three parameter combinations of $T=6,\gamma=1$; $T=2,\gamma=2$ and $T=1.2,\gamma=4$ (the performance of the three parameter combinations for different functions is slightly different. We also tested other parameter combinations and found that no combination with a product of $T$ and $\gamma$ less than $4$ can achieve near machine precision).
@@
Fig. 12 shows the numerical results for the first case. As mentioned above, the number of nodes $M$ required for M1 and M2 is determined by the frequency of the boundary interval, but the re-solution ability of polynomial approximation is weaker than that of trigonometric approximation, so the value of $M$ required by M2 is larger than M1; while M3 and M4 are determined by the maximum frequency of the function in the entire interval, and the number of nodes required is slightly larger than M1. At this time, the number of nodes $M$ required by the four algorithms to achieve machine accuracy was relatively small, therefore, there was no obvious difference in performance.
@@
In the third case, because the function oscillates near the boundary, the algorithm in this paper cannot avoid this singularity. For functions $f_{7}$ and $f_{8}$, the maximum frequency of the function appears in the boundary interval. According to (27), the number of nodes required by the proposed algorithm is $T_{\Delta}\gamma_{\Delta}=6$ times the maximum frequency, whereas the number of nodes required by M3 and M4 is $T\gamma=4$ times the maximum frequency. This is consistent with the results presented in Figures 2 and 3. For $f_{9}$, because the internal frequency is higher, the number of nodes required by the proposed algorithm is less than that of M3 and M4. As for M2, we can see that its approximation ability for the boundary interval data is obviously weaker than that of M1, therefore, more nodes are needed to achieve near-machine accuracy.
We observe that in certain cases (e.g., Fig. 12(f)), Algorithm M2 exhibits superior accuracy over M1 for small node counts $M$. Since M2 employs fewer boundary nodes, its captured boundary intervals inherently exhibit smoother relative behavior compared to M1 under the same conditions. Notably, reducing the number of boundary nodes in M1 can similarly enhance early-stage accuracy; however, as demonstrated in prior tests (Section 5.2), such parameter configurations compromise M1’s ability to achieve near-machine precision with sufficient nodes.
@@ 6.2 Algorithm improvement: boundary grid refinement
$L_{\Delta}^{R}=2\times\left\lceil T_{\Delta}\times(m^{R}_{\Delta}-1)\right\rceil,$
$h^{R}=\frac{2\pi}{L^{R}_{\Delta}},\quad x_{j}^{R}=(j-1)h,\quad j=1,2,\ldots,L^{R}_{\Delta}.$
@@
$J^{R}_{\Delta,1}=\{1,2,\ldots,m^{R}_{\Delta}\},~J^{R}_{\Delta,2}=\{\frac{L^{R}_{\Delta}}{2}+1,\frac{L^{R}_{\Delta}}{2}+2,\ldots,\frac{L^{R}_{\Delta}}{2}+m^{R}_{\Delta}\},$~
$J^{R}_{\Delta}=J^{R}_{\Delta,1}\cup J^{R}_{\Delta,2}.$
@@
The extension function $g^{R}_{c}(x^{R}_{j}),j=1,2,\ldots,L^{R}_{\Delta}$ was also obtained according to the method described in section 3. And then we obtain a function $f^{R}_{c}$ with a period of $2+\lambda$, where $\lambda$ is same as in (9). The data of the extension function $f^{R}_{c}(t_{\ell}),\ell=-M,\ldots,M,M+1,\ldots,M+\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)-1$ in one period $1,1+\lambda)$ are
@@
(a) the approximation error against $R$
(b) the approximation error against $t$ $(\omega=300)$
Figure 15: The approximation error with various $R$ for functions $\exp(\text{i}\omega\pi t)$, $M=500$.
@@
Figure 16: Comparisons of algorithms M1 and M5.
Now we take $\gamma=1,T_{\Delta}=6,m_{\Delta}=25$ to test the influence of parameter $R$ on the algorithm. As can be seen from Fig. 15(a), when $R\geq 5$, the approximation error not only does not improve, but becomes worse, which is caused by the instability of the calculation of the extension part. When $R\leq 4$, it can be observed that the refinement of grid near the boundary improves the approximation error. When $\omega=50$, the improvement is not obvious, because the number of nodes $M=500>T_{\Delta}\gamma_{\Delta}\omega$ has exceeded the threshold for resolving this frequency function. Therefore, when facing high-frequency functions, we recommend using parameters
@@
for calculation. It can be seen from Fig. 15(b) demonstrates that progressive increases in parameter $R$ from $1$ to $4$ induce uniform error reduction across the entire computational domain. As the boundary grid is refined, the reduction of the boundary error improves the overall approximation error.
@@
Next, we compare the performance of the M5 method with Chebyshev interpolation. Polynomial interpolation at Chebyshev nodes is a widely used approach for function approximation, renowned for its high accuracy and favorable resolution constant. Figure 17 shows the approximation errors for both methods applied to the test functions $f_{3}$, $f_{5}$, $f_{7}$, and $f_{9}$ as the number of interpolation nodes $M$ increases. The relative performance of the two methods depends primarily on the characteristics of the function, especially its frequency distribution properties. Overall, the performance difference between the methods is modest. Furthermore, both techniques are compatible with FFT-based implementations, ensuring high computational efficiency.
(a) $f_{3}(t)$
(b) $f_{5}(t)$
(c) $f_{7}(t)$
(d) $f_{9}(t)$
Figure 17: Comparisons of algorithms M5 and the Chebyshev interpolation.
@@ 7 Conclusions
Conflict of Interest declaration
The authors declare that they have no affiliations with or involvement in any organization or entity with any financial interest in the subject matter or materials discussed in this manuscript.
@@ Acknowledgments
The authors would like to sincerely thank Mark Lyon for generously sharing the MATLAB codes to facilitate comparative analysis in this work. We are also very grateful to the reviewers and the editor for their constructive comments.